A movie star, unwilling to give his age, posed the following riddle to a gossip columnist. " Seven years ago, I was eleven times as old as my daughter. Now I am 4 times as old as she is." How old is the star?
Answer by ankor@dixie-net.com(22740)

let s = star's present age
let d = daughter's present age
:
Write an equation for each statement:
:
"Seven years ago, I was eleven times as old as my daughter."
s - 7 = 11(d-7)
s - 7 = 11d - 77
s = 11d - 77 + 7
s = 11d - 70
:
"Now I am 4 times as old as she is."
s = 4d
:
How old is the star?
:
Find the daughters age first
replace s with 4d in the first simplified equation
4d = 11d - 70
4d - 11d = -70
-7d = -70
d = 10 yrs is the daughters age
then
4(10) = 40 yrs is the stars age
:
:
Check solutions in the first statement
40 - 7 = 11(10 - 7)